OpenStudy (anonymous):

Five cards are dealt from a standard 52-card deck. What is the probability that we draw a 1 ace, 1 two, 1 three, 1 four, and 1 five (this is one way to get a “straight”)?

OpenStudy (anonymous):

four choices for each up top right? and \[\binom{52}{5}\] for the denominator

OpenStudy (anonymous):

Yes, I've got the denom!

OpenStudy (anonymous):

The probability of each on their own is 4/52, but I know that changes as you draw each down to 4/51, 4/50, etc...but how do I express that?

OpenStudy (anonymous):

I still don't know what goes in the denominator. I can reliably determine the denominator, but I have no clue what's on top!

OpenStudy (anonymous):

the denomator is the number of ways you can choose 5 cards from a set of 52

OpenStudy (anonymous):

OH sorry, I meant numerator!!

OpenStudy (anonymous):

in other words "52 choose 5" written as \[\binom{52}{5}\] or sometimes \[_{52}C_5\]

OpenStudy (anonymous):

I meant I don't know what goes in the numerator! Silly me! Typo.

OpenStudy (anonymous):

computed via \[\frac{52\times 51\times 50\times 49\times 48}{5!}\]

OpenStudy (anonymous):

I was thinking, yes yes I know the denominator. Why is he/she telling me? I realize my mistake. Numerator is the concern.

OpenStudy (anonymous):

ooh numerator

OpenStudy (anonymous):

Yeah sorry! Thanks for your patience!

OpenStudy (anonymous):

4 choices for each, so \(4\times 4\times 4\times 4\times 4\)

OpenStudy (anonymous):

I don't really get where that's coming from.. hmm. Why can it just be 4s?

OpenStudy (anonymous):

4 choices of aces right?

OpenStudy (anonymous):

Yes, but there's more going on it seems. Why is it not including the reducing total or something like 4/52, 4/51, 4/50, /4/49

OpenStudy (anonymous):

that is taken care of in the denominator

OpenStudy (anonymous):

It is? Ahh alright...Your answer is right, but I'll have to ask my teacher for clarification about why it's taken care of. It's late and I won't mess with that. Thank you!

OpenStudy (anonymous):

think of it this way it is all the "counting principle" if the number of ways to do one thing is n and another m, then there are \(n\times m\) ways to go it together

OpenStudy (anonymous):

4 ways to get an ace, 4 ways to get a 2 etc so \(4^5\) ways all together

OpenStudy (anonymous):

Okay, I think I understand. The numerator is always the wild card for me

OpenStudy (anonymous):

then the number of ways to pick the 5 cards is 52 choose 5, which is again really the counting principle in action

OpenStudy (anonymous):

if it was for example 3 spades and two clubs, the answer would be \[\frac{\binom{13}{3}\binom{13}{2}}{\binom{52}{5}}\]

OpenStudy (anonymous):

these kind of problems, the question is always "how many ways"

OpenStudy (anonymous):

I don't get what you've said. I'm sorry.

OpenStudy (anonymous):

I get how they apply in their own situations, but I'm SO bad at generalizing these sorts of rationales

OpenStudy (anonymous):

The next question is asking me how ANY straight could be obtained and I don't even know where to start.

OpenStudy (anonymous):

oh that is ok we are almost there

OpenStudy (anonymous):

Like what goes on top? No idea.

OpenStudy (anonymous):

it's no longer just 4s, and it's not comb(13,2) either. So it's a complete effing mystery to my mind.

OpenStudy (anonymous):

we got one straight right? one that begins with an ace and ends with a 5

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

ok so now we know the probability of that particular straight (that is why that question came first)

OpenStudy (anonymous):

now all we have to do is figure out how many different possible straights there are

OpenStudy (anonymous):

no idea!

OpenStudy (anonymous):

so we count them!

OpenStudy (anonymous):

not a poker player i guess huh?

OpenStudy (anonymous):

Not in a million years.

OpenStudy (anonymous):

ace to 5 2 to 6 3 to 7 4 to 8 5 to 9 6 to 10 7 to jack 8 to queen 9 to king 10 to ace

OpenStudy (anonymous):

i count ten

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

btw you can't have for example jack to 2

OpenStudy (anonymous):

so we are done

OpenStudy (anonymous):

multiply the above answer by 10

OpenStudy (anonymous):

oh. So .000394 *10

OpenStudy (anonymous):

i.e. \[\frac{10\times 4^5}{\binom{52}{5}}\]

OpenStudy (anonymous):

oh if you had the previous answer then yeah move the decimal one to the left

OpenStudy (anonymous):

right

OpenStudy (anonymous):

whatever

OpenStudy (anonymous):

That's not quite right.

OpenStudy (anonymous):

The answer is .00355

OpenStudy (anonymous):

i get \[0.00394\]

OpenStudy (anonymous):

So did I, from what you told me. The textbook says it's .00355

OpenStudy (anonymous):

they are wrong

OpenStudy (anonymous):

umm..

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

unless...

OpenStudy (anonymous):

you are supposed to subtract the number of flushes which i doubt

OpenStudy (anonymous):

I have no idea. This is all bullcrap to me. I don't like math.

OpenStudy (anonymous):

Let STATA do it, that's what I say.

OpenStudy (anonymous):

But I'll put down the answer we worked out and ask later.

OpenStudy (anonymous):

Thanks for your help.